### v1

# MultiGran-STGCNFog: Towards Accurate and High-Throughput Inference for Multi-Granular Spatiotemporal Traffic Forecasting

Abstract
Accurate traffic forecasting and swift inference provision are essential for intelligent transportation systems. However, the present Graph Convolutional Network (GCN)-based approaches cannot extract and fuse multi-granular spatiotemporal features across various spatial and temporal scales sufficiently, proven to yield less accurate forecasts. Besides, additional feature extraction branches introduced in prior studies critically increased model complexity and extended inference time, making it challenging to provide fast inference for traffic forecasting. In this paper, we propose MultiGran-STGCNFog, an efficient fog distributed inference system with a novel traffic forecasting model that employs multi-granular spatiotemporal feature fusion on generated dynamic traffic graphs to fully capture interdependent traffic dynamics. The proposed scheduling algorithm GA-DPHDS, optimizing layer execution order and layer-device scheduling scheme simultaneously, contributes to considerable inference throughput improvement by leveraging heterogeneous fog devices in a pipelined manner. Extensive experiments on real-world datasets demonstrate the superiority of the proposed method over selected baselines.
1 Introduction
With the advancement of Internet of Things (IoT) technology, making cities smarter through IoT has become a prominent research trend. The large volume of data collected by IoT devices has shifted attention toward data analysis algorithms and corresponding servings to enhance transportation system intelligence Zantalis et al. (2019). Traditional centralized transportation systems are limited by high latency and weak scalability. Against this backdrop, fog computing, which processes data on nearer nodes close to the data source, has emerged as a more efficient and scalable solution.
Having IoT sensors set up on roads and a large volume of traffic data gathered, such data needs to be analyzed efficiently to predict future traffic conditions. Accurate predictions benefit services like transportation optimization, navigation planning, and congestion prevention Li et al. (2022). Consequently, traffic forecasting has become an essential task, focusing on predicting road network patterns using historical observations Liu et al. (2023).
While GCN-based methods dominate spatiotemporal forecasting, transformer-based models have recently shown advantages in capturing long-term dependencies Jiang et al. (2023); Liu et al. (2023). However, GCNs still remain irreplaceably unmatched in handling spatial dependencies, and with carefully designed architectures, they are capable of achieving better performance than transformers Dan et al. (2024); Zhang et al. (2024). Considering their advanced technological maturity and broad applicability in spatiotemporal traffic-related tasks, we focuse on GCN-based approaches.
Existing GCN and transformer models primarily design complex Spatial-Temporal (S-T) blocks but lack a structured approach to extract and integrate hierarchical, multi-granular spatiotemporal features across spatial and temporal dimensions. Although a few works have explored multi-granular spatial Wang et al. (2021); Guo et al. (2021) or temporal features Wang and Jing (2022); Huo et al. (2023); Guo et al. (2019) independently, seamlessly integrating both is critical for accurate forecasting and remains unexplored.
Previous studies manipulating multi-scale traffic data introduced multiple feature extraction branches to capture features of different granularities. This significantly increased model complexity, leading to more extended inference time, which is challenging for Graph Neural Network (GNN) serving where efficiency and low latency are critical. No solutions were proposed in these earlier works to address this issue.
Besides, designing accurate traffic forecasting models with swift inference still encounters several key challenges:
1) Traffic data exhibits highly dynamic patterns, making it difficult to model spatiotemporal correlations.
2) In GNN serving systems, efficiently leveraging available resources is essential for speeding up where communication latency becomes a major hurdle Zeng et al. (2022).
3) Optimizing distributed parallel inference schemes is increasingly complex with the exponentially-growing search space and layer input-output dependencies.
4) Mainstream distributed parallel frameworks, such as GPipe Huang et al. (2019), assume even partitioning of stages, ignoring hardware heterogeneity. Real-word device clusters are often heterogeneous, and model layers cannot be fully in even distribution to ensure equal processing times.
To fill these gaps, we propose MultiGran-STGCNFog, an efficient GNN inference system that extracts and fuses multi-granular spatiotemporal features across various spatial and temporal scales. The distributed pipeline-parallel architecture across fog nodes enables MultiGran-STGCNFog to tackle the extended inference latency. These designs support swift traffic forecasting while enhancing forecasting performance.
Our contributions can be summarized as follows:
  • A novel GCN, titled Multi-Granularity Spatiotemporal Graph Convolution Network (MultiGran-STGCN), is developed. It employs dynamic graph techniques and a multi-granular spatiotemporal feature fusion mechanism through Laplacian-driven hierarchical graph clustering and multi-scale temporal modeling.
  • To mitigate the prolonged inference latency induced by additional feature extraction branches in multi-scale modeling, we construct MultiGran-STGCNFog to accelerate inference, leveraging the resources of heterogeneous fog devices. It enables faster inference by pipeline parallelism, which allows immediate data transmission among fog nodes in proximity to avoid remote transfers.
  • We develop a robust heterogeneous cross-device scheduling algorithm, GA-DPHDS, which jointly optimizes the layer execution order and the layer-device scheduling strategy, overcoming the search space explosion challenge. Our experiments demonstrate the necessity of layer execution order optimization, which is overlooked in previous studies with pipeline parallelism.
  • The proposed framework is tested with physical computing nodes and evaluated using three real-world traffic datasets. Experimental results demonstrate its superior performance compared to selected baselines.
2 Related Work
2.1 GCNs: Spatiotemporal Traffic Forecasting
When combined with temporal modeling techniques like RNNs, GCN-based models are particularly effective at capturing spatiotemporal dependencies Seo et al. (2018); Cui et al. (2020). DCRNN Li et al. (2017) modeled traffic as a diffusion process on directed graphs, using bidirectional diffusion convolution and RNNs to outperform baselines. Gated-STGCN Yu et al. (2017) leveraged graph convolutions for spatial and gated 1D convolutions for temporal modeling, improving efficiency by removing recurrent structures. ASTGCN Guo et al. (2019) employed attention mechanisms to capture spatial and temporal dependencies, while HGCN Guo et al. (2021) modeled traffic across micro- and macro-level networks. GWNET Wu et al. (2019) combined dilated causal and graph convolutions for hidden dependency capture, achieving state-of-the-art results, and OGCRNN Guo et al. (2020) integrated GCN with recurrent units for enhanced feature extraction.
2.2 Distributed and Parallel GNNs
The growing computational demands of deep learning and the inability to deliver fast serving have driven a shift from centralized to distributed computing, leveraging parallelism in GNNs. Data parallelism partitions input data across nodes, using graph parallelism to minimize inter-node communication or mini-batch parallelism with node sampling. Model parallelism distributes the model itself through operator parallelism (simultaneous operations), pipeline parallelism (processing stages sequentially), or ANN parallelism (spreading layers across devices) Besta and Hoefler (2024).
GPipe Huang et al. (2019) applied pipeline parallelism by dividing models into stages across accelerators for micro-batch processing, while PipeDream Narayanan et al. (2019) enhanced this with the 1F1B algorithm for concurrent forward and backward passes. A recent work, GNNPipe Chen et al. (2023) optimized GNN layer distribution across GPUs, reducing communication overhead and enabling hybrid parallelism for large-scale graphs.
3 Problem Formulation
In this section, we formally model the traffic forecasting task within the framework of graph-based representation and distributed pipeline-parallel inference to define system objectives, laying the foundation for our proposed framework.
3.1 Traffic Forecasting
We model the traffic network as a graph $G=(V,E,A)$, where $V$ is the set of traffic observation points with $N$ nodes, i.e., $|V|=N$. $A\in\mathbb{R}^{N\times N}$ defines the adjacency and $E$ is the set of edges, where $e_{ij}\in E$ denotes a connection between nodes $v_{i}$ and $v_{j}$. Each node $v_{i}$ at time $t$ has a $d$-dimensional feature vector $x_{i}^{t}\in\mathbb{R}^{d}$. The entire traffic network at time $t$ is represented by the feature matrix $X_{t}\in\mathbb{R}^{N\times d}$, with each row corresponding to the feature vector of a node.
The traffic forecasting task aims to predict the future traffic state $\mathcal{X}_{t+1:t+T^{\prime}}=[X_{t+1},\ldots,X_{t+T^{\prime}}]$ of $T^{\prime}$ future time steps from $t+1$ to $t+T^{\prime}$, based on historical traffic data $\mathcal{X}_{t-T+1:t}=[X_{t-T+1},\ldots,X_{t}]$ of $T$ time steps from $t-T+1$ to $t$. The objective is to learn a nonlinear mapping $f(\cdot)$, which captures the spatiotemporal dependencies from historical traffic observations to predict future traffic states:
$\mathcal{X}_{t-T+1:t}=[X_{t-T+1},\ldots,X_{t}]\xrightarrow{f(\cdot)}[X_{t+1},\ldots,X_{t+T^{\prime}}].$ (1)
3.2 Pipeline-Parallel Model Inference
Given a GNN model $\mathcal{M}$ for traffic forecasting, it consists of $L$ layers, denoted as $l_{j}$ (where $j=1,2,\dots,L$). Each layer $l_{j}$ is characterized by its memory consumption $\text{mem}_{j}$, input parameter size $\text{input}_{j}$, and output parameter size $\text{output}_{j}$. To facilitate a more streamlined inference process, we partition the model $\mathcal{M}$ into multiple stages and subdivide a mini-batch of size $B$ into multiple micro-batches, each of size $B_{\mu}$, allowing for better resource utilization and overlap of computation and communication. Each stage $\text{stage}_{k}$ is defined as:
$\text{stage}_{k}=\{l_{start_{k}},l_{end_{k}}\},\quad\bigcup_{k=1}^{K}\text{stage}_{k}=\mathcal{M}.$ (2)
Every stage consists of a set of contiguous model layers $l_{1},l_{2},\dots,l_{L}$, and is assigned to a single device. This process requires solving a combinatorial optimization problem to balance performance and resource constraints. Generated stages are allocated to a heterogeneous fog cluster that has the device set $\mathcal{D}=\{d_{1},d_{2},\dots,d_{Nd}\}$ containing $N_{d}$ heterogeneous devices, each with distinct computational capacity, memory capacity, and communication bandwidth. The layer-device scheduling strategy $S$ is defined as the mapping:
$S$ =$=\{(d_{i},\text{stage}_{k})\mid i,k\in\{1,2,\dots,n\},\;n\leq N_{d};$ (3)
$\quad\quad\quad\quad\quad\quad L_{k}=\{l_{start_{k}},\dots,l_{end_{k}}\}\}.$
Our goal is to find a joint strategy $(\mathcal{O},S)$, with the layer execution order $\mathcal{O}$ and the layer-device scheduling strategy $S$, that determines the partitioning of all layers of $\mathcal{M}$ into stages and their allocation to heterogeneous fog devices for conducting distributed parallel inference in a pipelined manner. $(\mathcal{O},S)$ leads to the two optimization objectives below.
3.2.1 Minimizing Longest Stage Execution Time
In pipeline parallelism, maximizing pipeline throughput is equivalent to minimizing the execution time of the longest stage. The execution time $T_{\text{exec}}(d_{i})$ for device $d_{i}$ involves computation time $T_{\text{comp}}(\text{stage}_{i},d_{i})$ and communication time $T_{\text{comm}}(\text{stage}_{i},d_{i-1},d_{i})$ for intermediate value transmission:
$T_{\text{exec}}(d_{i})=\max\left(\begin{aligned} &T_{\text{comp}}(\text{stage}_{i},d_{i}),\\ &T_{\text{comm}}(\text{stage}_{i},d_{i-1},d_{i})\end{aligned}\right),$ (4)
$T_{\text{comp}}(\text{stage}_{i},d_{i})=\sum_{l_{j}\in\text{stage}_{i}}\text{proc}_{i}(l_{j}),$ (5)
$T_{\text{comm}}(\text{stage}_{i},d_{i-1},d_{i})=\frac{\text{output}_{l_{end_{i}}}\cdot B_{\mu}}{\min(b_{i-1}^{\text{up}},b_{i}^{\text{down}})}.$ (6)
$\text{Proc}_{i}(l_{j})$ represents the processing time for layer $l_{j}$ on device $d_{i}$, and the available communication bandwidth between two devices, $b_{i-1,i}$, is bounded by the sender’s up-link bandwidth $b_{i-1}^{\text{up}}$and the receiver’s down-link bandwidth $b_{i}^{\text{down}}$.
3.2.2 Load Balancing
To mitigate pipeline inefficiencies caused by load imbalance, We aim to balance the load across devices by minimizing the standard deviation $\sigma$ of devices’ execution time $T_{\text{exec}}(d_{i})$:
$\sigma=\sqrt{\frac{1}{N_{d}}\sum_{i=1}^{N_{d}}\left(T_{\text{exec}}(d_{i})-\overline{T_{\text{exec}}}\right)^{2}}.$ (7)
3.2.3 Optimization Summary
In conclusion, to maximize the pipeline throughput, we optimize both the layer execution order $\mathcal{O}$ and the layer-device scheduling strategy $S$, searching for the optimal $(\mathcal{O}^{*},S^{*})$ that minimizes the maximum execution time across all devices:
$\min_{(\mathcal{O}^{*},S^{*})}\max_{1\leq i\leq N_{d}}\left\{\begin{aligned} \max\big{(}&T_{\text{comp}}(\text{stage}_{i},d_{i}),\\ &T_{\text{comm}}(\text{stage}_{i},d_{i-1},d_{i})\big{)}\end{aligned}\right\}.$ (8)
Figure 1: The network architecture of MultiGran-STGCN.
4 System Design: MultiGran-STGCNFog
This section begins with a system overview illustrating all phases of the proposed inference system. Next, MultiGran-STGCN components are elaborated in detail, along with the heterogeneous cross-device execution scheduling algorithm.
4.1 System overview
The workflow of MultiGran-STGCNFog consists of four phases: (1) Registration Phase: Fog nodes connect to the fog manager for node registration and initialization. (2) Profiling Phase: The fog manager distributes the forecasting model $\mathcal{M}$ and sampled historical traffic data to registered nodes, where inference is performed to profile computation time $\text{proc}_{i}(l_{j})$, memory usage $\text{mem}_{j}$, input and output parameter sizes, $\text{input}_{j}$ and $\text{output}_{j}$ for every layer, and P2P communication bandwidth $b_{i,i+1}$. (3) Scheduling Phase: The fog manager calculates the optimal scheme $(\mathcal{O}^{*},S^{*})$ and assigns stage allocation with corresponding model segments. (4) Runtime Phase: Fog nodes perform inference collaboratively, exchanging intermediate values as needed.
4.2 Traffic forecasting model
The architecture of MultiGran-STGCN is depicted in Fig. 1, and its four components will be introduced sequentially.
4.2.1 Multi-scale Traffic Data Modeling
Three spatial scales are constructed for multi-spatial granularity generation. With the distance matrix $D$, distances are converted into a similarity measure via a Gaussian kernel:
$A(i,j)=\exp\left(-\frac{d_{ij}}{\text{var}(d_{i})}\right),\quad 0\leq A(i,j)\leq 1.$ (9)
The Laplacian matrix $L$ is computed and normalized as:
$L=D_{\text{deg}}-A,\quad L_{\text{norm}}=D_{\text{deg}}^{-\frac{1}{2}}LD_{\text{deg}}^{-\frac{1}{2}}.$ (10)
Clustering is applied based on the eigenvectors H, which is gained by eigenvalue decomposition on $L_{\text{norm}}$, minimizing:
$\min\sum_{i=1}^{n}\sum_{j=1}^{k}\left\|\textbf{H}_{i}-\mu_{j}\right\|^{2},$ (11)
where $\mu_{j}$ is the centroid of cluster $j$. Each cluster is treated as a node, $\mathbf{X}_{i}\in\mathbb{R}^{T\times d}$, in the coarser spatial scale, and features are aggregated through pooling operations:
$\mathbf{X}_{\text{min}}^{(j)}$ $=\min_{i\in\mathcal{C}_{j}}\mathbf{X}_{i},\;\mathbf{X}_{\text{mean}}^{(j)}=\frac{1}{|\mathcal{C}_{j}|}\sum_{i\in\mathcal{C}_{j}}\mathbf{X}_{i},\;\mathbf{X}_{\text{max}}^{(j)}=\max_{i\in\mathcal{C}_{j}}\mathbf{X}_{i},$ (12)
$\quad\mathbf{X}_{\text{cluster}}^{(j)}=\text{concat}\left(\mathbf{X}_{\text{min}}^{(j)},\mathbf{X}_{\text{mean}}^{(j)},\mathbf{X}_{\text{max}}^{(j)}\right).$
Three different temporal scales are set up based on original traffic data $\mathbf{X}\in\mathbb{R}^{T\times N\times d}$, namely short-term, medium-term, and long-term scales, corresponding to hourly data $\mathbf{X}_{\text{hour}}\in\mathbb{R}^{T_{h}\times N\times d}$, daily data $\mathbf{X}_{\text{day}}\in\mathbb{R}^{T_{d}\times N\times d}$, and weekly data $\mathbf{X}_{\text{week}}\in\mathbb{R}^{T_{w}\times N\times d}$. In detail, short-term modeling captures local dynamic changes over recent time steps. Setting the short-term window length as $W_{s}$, traffic data from $t_{0}-h+1$ to $t_{0}$ is extracted, to construct the short-term scale $\mathbf{S}_{t}=\left[\mathbf{X}_{t_{0}-h+1},\mathbf{X}_{t_{0}-h+2},\dots,\mathbf{X}_{t_{0}}\right]\in\mathbb{R}^{W_{s}\times N\times d}$, using the short-term offset $h$. Mid-term scale captures periodic traffic patterns at daily peaks, formulated as $\mathbf{M}_{t}=\left[\mathbf{X}_{t_{0}-\delta_{d}\cdot q},\mathbf{X}_{t_{0}-\delta_{d}\cdot(q-1)},\dots,\mathbf{X}_{t_{0}-\delta_{d}}\right]\in\mathbb{R}^{W_{d}\times N\times d}$, where $\delta_{d}$ is the time offset for daily periodicity, $W_{d}$ is the mid-term window length. Similarly, the long-term scale $\mathbf{L}_{t}=\left[\mathbf{X}_{t_{0}-7\cdot\delta_{d}\cdot q},\mathbf{X}_{t_{0}-7\cdot\delta_{d}\cdot(q-1)},\dots,\mathbf{X}_{t_{0}-7\cdot\delta_{d}}\right]\in\mathbb{R}^{W_{w}\times N\times d}$, models trend changes over larger time scales, capturing weekly pattern from the same time points. Afterward, each temporal scale is trimmed according to the minimum number of samples across the three scales $T_{\text{min}}=\min(T_{h},T_{d},T_{w})$.
4.2.2 Dynamic Graph Generation Block
Unlike the traditional graph convolution operator, defined by a static graph structure $\hat{A}$:
$H^{(l+1)}=\sigma\left(\hat{D}^{-\frac{1}{2}}\hat{A}\hat{D}^{-\frac{1}{2}}H^{(l)}W^{(l)}\right).$ (13)
It is designed to dynamically generate the adjacency matrix $A_{t}$ at each time step, allowing the graph structure to adapt to changing input data. For each node $i$, two embedding vectors $\mathbf{v}_{1,i}\in\mathbb{R}^{d}$ and $\mathbf{v}_{2,i}\in\mathbb{R}^{d}$ are assigned, representing the roles of nodes as senders and receivers during the graph information dissemination. Dynamic adjacency matrix $A_{t}$ is generated through node similarity calculation by the inner product:
$A_{ij}=\text{ReLU}(\mathbf{v}_{1,i}\cdot\mathbf{v}_{2,j})=\text{ReLU}\left(\sum_{k=1}^{d}v_{1,i,k}\cdot v_{2,j,k}\right).$ (14)
4.2.3 Feature Extraction Branches
We introduced a Cross-Granularity Dynamic Fusion Block (C-GF block) to further enhance the extraction of spatiotemporal correlations, which integrates feature across different scales by adjusting the fusion weights of features from different granularity levels dynamically.
The primary components of a S-T Block include Temporal Convolution, Graph Convolution, and Temporal Attention Mechanism. The temporal convolution extracts temporal dependencies from input data $X\in\mathbb{R}^{B\times C\times N\times T}$:
$X_{\text{temp}}=\tanh(W_{\text{time}}*X+b_{\text{time}}),$ (15)
where $W_{\text{time}}$ is the temporal convolution kernel. To enhance feature extraction, $X_{\text{temp}}$ is then split into two parts:
$X^{\prime}_{\text{temp}}=\tanh(X_{\text{temp}_{1}})\cdot\sigma(X_{\text{temp}_{2}}).$ (16)
The graph convolution captures spatial dependencies with the dynamically generated topology $A_{\text{dynamic}}$ by (14) :
$X_{\text{gcn}}=\tilde{D}^{-1/2}A_{\text{dynamic}}\tilde{D}^{-1/2}X^{\prime}_{\text{temp}}W_{\text{gcn}}.$ (17)
Having $A_{\text{dynamic}}$ representing weighted spatial connections between nodes, it is similar to the temporal attention mechanism that dynamically assigns weights to different time steps, leveraging Scaled Dot-Product Attention with query (Q), key (K), and value (V) matrices:
$\begin{split}Q=X_{\text{gcn}}W_{Q},\quad K=X_{\text{gcn}}W_{K},\quad V=X_{\text{gcn}}W_{V},\\ \text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V.\end{split}$ (18)
C-GF blocks integrate multi-granular feature representations through learnable gates that regulate the contribution of each input feature set to the final fused representation. Features from spatial scales: $X_{L1}$, $X_{L2}$, and $X_{L3}$ are first concatenated to form a joint representation $X_{\text{concat}}$. Subsequently, the gating mechanism adjusts the contribution by the learnable weight matrix $W_{g}$ followed by a fusion step:
$\mathbf{G}=\sigma(W_{g}\cdot X_{\text{concat}}+b_{g}),$ (19)
$X_{\text{fused}}=\mathbf{G}_{L1}\odot X_{L1}+\mathbf{G}_{L2}\odot X_{L2}+\mathbf{G}_{L3}\odot X_{L3},$ (20)
$X_{\text{fusion}}=\sigma(W_{\text{fusion}}\cdot X_{\text{fused}}+b_{\text{fusion}}).$ (21)
4.2.4 Traffic Forecasting Head
In the forecasting head, to enable dynamical adjustment on the contribution of each time scale based on their importance, learnable weight matrices are exploited to compute the final prediction operating the output features $x_{\text{hour}}$, $x_{\text{day}}$, and $x_{\text{week}}$:
$x=W_{\text{hour}}\odot x_{\text{hour}}+W_{\text{day}}\odot x_{\text{day}}+W_{\text{week}}\odot x_{\text{week}}.$ (22)
4.3 Cross-device execution scheduling
Explained by (8), $(O^{*},S^{*})$ is crucial in determining how layers should be orchestrated and assigned. It is worth noting that previous works often ignored the optimization for the layer execution order $O$. However, our experiments show that the layer execution order plays a decisive role regarding the pipeline inference throughput. We model the input-output dependencies between layers using a Directed Acyclic Graph (DAG), where each node represents a layer $l_{j}$. If the output of layer $l_{j}$ is required as input for layer $l_{j+1}$, a directed edge is established from $l_{j}$ to $l_{j+1}$. This DAG structure captures the sequential and parallel execution constraints between layers. As the number of layers $L$ increases, the search space for both $O^{*}$ and $S^{*}$ grows exponentially. Hence, we propose GA-DPHDS, an algorithm utilizing the NSGA-II genetic algorithm with dynamic programming for optimization.
First, the population Pop is initialized randomly, where each individual $\text{ind}_{i}$ represents a layer execution order $O$, and individuals must satisfy the topological ordering constraints imposed by the DAG. For each $\text{ind}_{i}$, the overall inference time $T_{\text{overall}}(\text{ind}_{i})$, which is the longest device execution time; and the load balancing factor $\sigma(\text{ind}_{i})$, the standard deviation of execution times across all devices, are calculated:
$T_{\text{overall}}=\max_{d_{i}\in\mathcal{D}}\max\big{(}$ $\sum_{l_{j}\in\text{stage}_{i}}T_{\text{comp}}(l_{j},d_{i}),$ (23)
$T_{\text{comm}}(d_{i-1},d_{i})\big{)},$
$\sigma(\text{ind}_{i})=\sqrt{\frac{1}{N_{d}}\sum_{i=1}^{N_{d}}\left(T_{\text{exec}}(d_{i})-\overline{T_{\text{exec}}}\right)^{2}}.$ (24)
Next, Fast Non-dominated Sorting is applied to rank the individuals based on their performance in both objective functions. We calculate the crowding distance of $\text{ind}_{i}$, which quantifies how isolated an individual is in the objective space:
$\text{Distance}(\text{ind}_{i})=$ $\,\frac{T_{\text{overall}}(\text{ind}_{i+1})-T_{\text{overall}}(\text{ind}_{i-1})}{T_{\text{overall, max}}-T_{\text{overall, min}}}$ (25)
$+\frac{\sigma(\text{ind}_{i+1})-\sigma(\text{ind}_{i-1})}{\sigma_{\text{max}}-\sigma_{\text{min}}}.$
Subsequently, selection, crossover, and mutation operations are performed. Tournament selection is based on rank and crowding distance, and crossover generates offspring under DAG constraints by exchanging segments of the parent individuals’ genes. Swap mutation introduces random changes to the execution order, still adhering to DAG constraints.
For each individual $\text{ind}_{i}$, we leverage dynamic programming to further optimize the layer-device scheduling strategy $S$, defining a two-dimensional dynamic programming table $T$, where $T[i][j]$ represents the minimal maximum execution time when assigning the first $i$ layers to the first $j$ devices. The state transition recurrence is given by:
$T[i][j]=\min_{k<i}\left(\max\left(T[k][j-1],\right.\right.$ (26)
$\left.\left.\qquad\qquad\quad T_{\text{stage}}(k+1,i,D_{j})\right)\right).$
This formula indicates that we need to find a splitting point $k$ such that the layers from $k+1$ to $i$ are assigned to the current device $D_{j}$, while the first $k$ layers have already been optimally assigned to devices $D_{1}$ to $D_{j-1}$. The minimal maximum execution time $T[i][j]$ consists of the maximum of:
1) The minimal maximum execution time of the first $k$ layers on first $j-1$ devices, $T[k][j-1]$.
2) Stage execution time on device $D_{j}$, given by
$T_{\text{stage}}(k+1,\ i,\ D_{j})=\max\left\{\begin{aligned} &T_{\text{comp}}(k+1,\ i,\ D_{j}),\\ &T_{\text{comm}}(D_{j-1},\ D_{j})\end{aligned}\right\}.$ (27)
Here, $T_{\text{comp}}(k+1,\ i,\ D_{j})$ represents the computation time for device $D_{j}$ to process layers $k+1$ to $i$, calculated as:
$T_{\text{comp}}(k+1,\ i,\ D_{j})=\sum_{l=k+1}^{i}T_{\text{comp}}(l,\ D_{j}),$ (28)
and $T_{\text{comm}}(D_{j-1},\ D_{j})$ is the communication time between devices $D_{j-1}$ and $D_{j}$, formulated as:
$T_{\text{comm}}(D_{j-1},D_{j})=\begin{cases}\frac{\text{output}_{k}\cdot B_{\mu}}{b_{j-1,j}},&\text{if }D_{j-1}\neq D_{j},\\ 0,&\text{if }D_{j-1}=D_{j}.\end{cases}$ (29)
With such a state transition, the dynamic programming table $T$ is updated iteratively to reflect the optimal solution $(O^{*},S^{*})$ to the cross-device execution scheduling.
5 Experiments
The experiments section aims to: (1) Validate the performance of MultiGran-STGCN. (2) Demonstrate the necessity of GA-DPHDS against homogeneous counterparts. (3) Explore the impact of device network bandwidth on overall inference throughput and identify system bottlenecks. (4) Evaluate the effectiveness of our optimization strategies.
Table 1: Heterogeneous device table
Device CPU Avail. Mem. (GB) Bandwidth (Gbps)
A i7-12700F @ 2.10 GHz 32 3.2
B Xeon E3-1230 v6 @ 3.50 GHz 16 2.4
C i7-9700K @ 3.60 GHz 48 1.6
D M3 Pro (11-core) 18 2.4
E i7-7700 @ 3.60 GHz 16 2.0
F i7-9750H @ 2.60 GHz 8 1.2
5.1 Experimental Setup
Table 2: Heterogeneous clusters configuration
Device Cluster
One Two Three Four Five Six Seven Eight Nine Ten
Device A
Device B
Device C
Device D
Device E
Device F
We conducted experiments using six heterogeneous fog devices with varying capabilities regarding the CPU, memory, and communication bandwidth. Table 1 provides an overview of these devices, which were configured into ten clusters with distinct memberships shown in Table 2, where network communication was enabled through a symmetric local Ethernet. We evaluated the performance of MultiGran-STGCNFog regarding the forecasting accuracy and inference throughput.
Three benchmark datasets, PEMS04, PEMS07, and PEMS08 were used to evaluate MultiGran-STGCN. Forecasting performance was compared against a set of statistical and GCN approaches introduced in Section 2, including HA, LSTM Cui et al. (2018), GRU Agarap (2018), GCRN Seo et al. (2018), Gated-STGCN Yu et al. (2017), GWNET Wu et al. (2019), OGCRNN Guo et al. (2020), HGCN Guo et al. (2021), and ASTGCN Guo et al. (2019). Baseline parameters were configured as reported in their respective works.
We leveraged MAE, MAPE, and RMSE to measure forecasting performance, and inference efficiency was assessed by the pipeline throughput on PEMS04 (samples/second), with network latency randomly perturbed between 10 and 30 ms to better simulate real-world fog environments.
Table 3: Performance on PEMS04, PEMS07, and PEMS08 datasets. Pink/green/yellow marks the best/second-best/third-best performance.
Metric HA LSTM GRU GCRN Gated-STGCN GWNET OGCRNN HGCN ASTGCN MG-single MG-wms MG-wmt MultiGran-STGCN
PEMS04 MAE 28.38 20.26 20.18 20.78 20.36 18.12 19.76 18.29 20.35 19.00 18.96 18.33 18.11
15 mins MAPE 19.99% 13.87% 13.56% 15.78% 15.77% 12.58% 13.93% 13.02% 14.23% 13.74% 13.56% 12.83% 12.45%
RMSE 41.82 31.71 31.61 31.67 31.14 28.83 30.49 29.84 31.92 29.92 29.74 29.08 28.84
MAE 31.77 22.31 22.23 22.14 21.97 18.86 20.47 19.10 20.93 19.84 19.80 19.04 18.66
30 mins MAPE 22.65% 15.39% 14.94% 16.68% 16.79% 13.13% 14.42% 13.67% 14.57% 14.38% 14.15% 13.21% 12.83%
RMSE 46.49 34.54 34.46 33.65 33.44 29.93 31.55 30.06 33.00 31.11 30.90 30.16 29.82
MAE 38.51 26.41 26.33 24.98 25.17 20.06 21.74 20.62 21.87 21.39 20.89 20.19 19.55
60 mins MAPE 28.20% 18.65% 17.93% 18.68% 19.05% 13.98% 15.41% 14.86% 15.10% 15.47% 14.81% 13.92% 13.45%
RMSE 55.76 40.07 40.08 37.69 37.87 31.62 33.37 32.20 34.82 33.23 32.77 31.89 31.31
PEMS07 MAE 32.82 21.79 21.83 22.71 22.07 19.02 20.50 19.18 22.76 20.87 18.65 18.88 18.56
15 mins MAPE 15.04% 9.26% 9.23% 11.59% 11.57% 8.20% 9.10% 8.29% 9.96% 11.29% 8.11% 8.25% 7.99%
RMSE 47.97 33.84 33.76 34.16 32.84 30.14 31.85 30.25 35.42 31.46 29.72 29.83 29.82
MAE 37.03 24.42 24.50 24.51 24.14 20.32 21.54 20.46 23.89 22.33 19.67 20.14 19.52
30 mins MAPE 17.19% 10.39% 10.36% 12.17% 12.40% 8.67% 9.50% 8.77% 10.30% 12.05% 8.51% 8.73% 8.38%
RMSE 53.99 37.56 37.52 36.83 35.95 32.16 33.48 32.27 37.93 33.71 31.60 31.79 31.55
MAE 45.33 29.51 29.61 28.00 28.08 22.37 23.16 22.31 25.32 24.35 20.90 21.84 20.76
60 mins MAPE 21.56% 12.80% 12.67% 13.45% 14.12% 9.42% 10.20% 9.54% 10.78% 13.29% 9.03% 9.49% 8.95%
RMSE 65.74 44.59 44.42 41.75 41.45 35.12 35.19 35.06 41.12 36.58 34.00 34.47 33.81
PEMS08 MAE 23.11 16.13 16.08 16.85 16.97 14.11 16.04 14.37 16.53 15.14 14.80 14.28 14.03
15 mins MAPE 14.46% 10.02% 10.03% 11.94% 13.58% 9.47% 10.66% 9.20% 10.62% 10.43% 10.37% 9.24% 9.34%
RMSE 34.13 24.95 24.89 25.20 24.92 21.65 24.43 21.92 25.24 22.78 22.59 21.95 21.90
MAE 26.08 17.94 17.86 17.59 17.99 14.84 16.77 15.23 17.01 16.10 15.37 15.05 14.49
30 mins MAPE 16.36% 11.06% 11.11% 12.27% 14.14% 9.84% 11.00% 9.62% 10.89% 11.09% 10.80% 9.61% 9.55%
RMSE 38.31 27.76 27.66 26.49 26.80 22.94 25.68 23.35 26.27 24.34 23.73 23.32 22.90
MAE 32.00 21.46 21.38 19.12 20.14 16.02 18.17 16.69 17.61 17.49 16.20 16.28 15.12
60 mins MAPE 20.28% 13.25% 13.47% 13.07% 15.27% 10.41% 11.79% 10.38% 11.26% 12.13% 11.43% 10.24% 9.91%
RMSE 46.50 32.79 32.73 29.97 30.31 24.88 27.87 25.58 27.56 26.48 25.35 25.39 24.18
5.2 Spatiotemporal Traffic Forecasting
Table 3 presents the performance of MultiGran-STGCN, baselines, and ablation variants: MG-single (single spatial and temporal granularity), MG-wms (without multi-spatial granularity), and MG-wmt (without multi-temporal granularity). MultiGran-STGCN consistently outperforms baselines and ablation models across 15, 30, and 60-minute horizons.
For 15-minute forecasts, MultiGran-STGCN achieves MAE values of 18.11, 18.56, and 14.03 on PEMS04, PEMS07, and PEMS08, respectively, surpassing all baselines. It also delivers optimal or near-optimal MAPE and RMSE. For 30-minute forecasts, MultiGran-STGCN continues to lead, notably achieving the best performance across every metric on all datasets and fully surpassing any minor discrepancies observed in the short-term forecasts. This illustrates that integrating multi-granular spatiotemporal modeling and designed dynamic mechanisms further enhances MultiGran-STGCN’s capability to capture complex traffic dependencies. For 60-minute forecasts, The MAE values for PEMS04, PEMS07, and PEMS08 are 19.55, 20.76, and 15.12, suggesting more significant performance advances. Furthermore, the ablation variants -wms and -wmt also show notable superiority over baselines. On PEMS07, -wms and -wmt achieve MAE, MAPE, and RMSE of 20.90, 9.03%, and 34.00, as well as 21.84, 9.49%, and 34.47, outpacing the best-performing baseline’s metric values of GWNET.
5.2.1 Ablation Study of Multi-spatiotemporal Scale Modeling
Figure 2: Ablation study: performance across prediction horizons.
Fig. 2 provides a more intuitive performance comparison among ablation variants and the complete MultiGran-STGCN model denoted as -full, to explore the performance gain of multi-scale modeling and multi-granular feature fusion design. On all three datasets, the -full model consistently outperforms other variants, and its relatively modest increase in errors indicates its effectiveness in capturing long-term dependencies, as the prediction horizon extends from 15 to 60 minutes. The enormous gap between the -single variant and other models points out the limitation of relying solely on single-dimensional scales, drawbacks of previous studies. This observation suggests that while -wms and -wmt partially incorporate multi-spatiotemporal feature extraction, the lack of comprehensive integration restricts their forecasting accuracy. Besides, it is also clear that as the prediction horizon increases, the error trajectories begin to diverge. The growing divergence underscores our model’s superiority in capturing long-term correlations and handling error accumulation.
5.3 Distributed pipeline-parallel model inference
5.3.1 Heterogeneous vs. Homogeneous Scheduling
Figure 3: Heterogeneous vs. homogeneous scheduling algorithm.
The scheduling methods of GPipe Huang et al. (2019) and PipeDream Narayanan et al. (2019) were adopted as the baseline against GA-DPHDS in terms of pipeline throughput. GPipe utilized homogeneous accelerators, while PipeDream partitioned DNN operators based on a single GPU’s capabilities. Devices and pipeline sequences were randomized, and each cluster was tested ten times to obtain the average.
As shown in fig. 3, GA-DPHDS outperformed the baseline across all cluster configurations, achieving throughput from 14.03 samples/s (Cluster 8) to 67.14 samples/s (Cluster 4), while the baseline peaked at 33.72 samples/s and dropped to 12.51 samples/s (Cluster 10). In Cluster 4, GA-DPHDS reached 67.14 samples/s, more than doubling the baseline’s 30.68 samples/s, highlighting its efficiency in resource-intensive setups. In Cluster 2, GA-DPHDS improved throughput to 49.59 samples/s compared to the baseline’s 22.08 samples/s. However, in Cluster 8, having only devices B and C, GA-DPHDS exhibited only minor improvements due to comparable device capabilities, consistent with table 4. These findings underscore GA-DPHDS’s effectiveness in optimizing throughput across varied cluster setups.
Table 4: Single device throughput (samples/s)
Device A B C D E F
Throughput 18.35 6.64 6.78 17.19 8.16 11.73
5.3.2 Exploring the Impact of Device Network Bandwidth
Figure 4: Cluster throughput with varying bandwidth scaling factors.
We explored the impact of device bandwidth on pipeline throughput by scaling, where bandwidth settings presented in Table 1 took the scaling factor of 1.0 as the standard. Fig. 4 shows that insufficient bandwidth (e.g., scaling factor 0.025) constrained intermediate data transfer efficiency, creating a system bottleneck. We found that in such cases, GA-DPHDS tended to allocate layers to the device with the highest computing power to minimize transmission costs, reducing pipeline parallelism. For instance, in Clusters 1–4, all layers were assigned to device A, resulting in low pipeline throughput. As the scaling factor increased, the average throughput improved (black line), with clusters containing more homogeneous devices showing higher sensitivity to bandwidth, while clusters 8–10 were less affected. Once bandwidth became sufficient (e.g., scaling factor 1.5), it ceased to be a bottleneck, and device computational capacity (cpu) emerged as the new limiting factor. These results demonstrate the adaptability of GA-DPHDS under varying bandwidth constraints.
5.3.3 Optimization Ablation
Figure 5: Ablation study for layer execution order optimization.
We examined the impact of layer execution order $O$ on throughput, optimized via the NSGA-II genetic algorithm. Without this optimization, layer sequences were randomly arranged, adhering only to input-output dependencies. As shown in Fig. 5, optimized layer execution significantly improved throughput. In Cluster 3, GA-DPHDS achieved 58.17 samples/s, a 41.04% increase compared to 41.24 samples/s without NSGA-II, marking the highest observed improvement. Similarly, in Cluster 4, with all the fog devices, throughput was improved by 27.4%, demonstrating that optimized execution order enhances pipeline parallelism and system efficiency. These results highlight the criticality of execution order optimization in maximizing pipeline throughput.
6 Conclusion and Future Work
In this work, we propose MultiGran-STGCNFog, an efficient GNN inference system with a novel traffic forecasting model. MultiGran-STGCNFog extracts spatiotemporal features across various spatial and temporal scales, and the distributed pipeline-parallel architecture of it enables high-performance inference throughput leveraging heterogeneous fog devices. Specifically, the dynamic mechanism and multi-granular feature fusion strengthen its capability to capture long-term traffic dependencies. The proposed scheduling algorithm GA-DPHDS, brings significant throughput improvement. In the future, we might explore the efficiency of our multi-granular spatiotemporal data modeling and feature fusion approach on transformer-based traffic forecasting models, and develop delicate work-stealing strategies to further balance pipeline workload during the runtime phase.